OpenStudy (anonymous):

Find the area of the region bounded by hyperbola (25x^2-4y^2=100) and the line x=3

OpenStudy (amistre64):

25x^2-4y^2=100 /100 ----------------- x^2 y^2 ------ - ------ = 1 100/25 100/4 x^2 y^2 ------ - ------ = 1 4 25 

OpenStudy (amistre64):

alomst lol .... x = 2 in that not 4

OpenStudy (unklerhaukus):

integrate drawing

OpenStudy (amistre64):